# Archilles’ Heel in Semi-open LLMs: Hiding Bottom against Recovery Attacks

### v1

@@ B.5 Evaluation Benchmarks
Code HumanEval Pass@1 0 Chen et al. (2021)
MBPP Pass@1 1 Austin et al. (2021)
Math GSM8K Exact Match 8 Cobbe et al. (2021)
Medical MMLU_Medical Accuracy 5 Hendrycks et al. (2021)
Finance FPB F1 0 Wang et al. (2023a)
Law LegalBench Accuracy 0 Guha et al. (2023)
Alignment MT-Bench Score (GPT-4) Zheng et al. (2023)
Evaluation on Resilience. We follow the Llama-2 report Touvron et al. (2023) to evaluate the recovered model, including 16 benchmarks, which are categorized into 6 groups. Table 11 summarizes the functionality benchmarks used in our experiments, along with their test methods and performance metrics. Our model ranks choices in multiple-choice tasks and generates answers for open-ended generation tasks.
@@
Commonsense Reasoning PIQA Accuracy 0 Bisk et al. (2020)
Hellaswag Accuracy 0 Zellers et al. (2019)
Winogrande Accuracy 0 Sakaguchi et al. (2019)
ARC_easy Accuracy 0 Clark et al. (2018)
ARC_challenge Accuracy 0 Clark et al. (2018)
Reading Comprehension OpenBookQ Accuracy 0 Mihaylov et al. (2018)
LAMBADA Accuracy 0 Paperno et al. (2016)
BoolQ Accuracy 0 Clark et al. (2019)
SQuADv2 HasAns_EM 2 Rajpurkar et al. (2018)
SQuADv2 HasAns_F1 2 Rajpurkar et al. (2018)
World Knowledge NaturalQuestions Exact Match 5 Kwiatkowski et al. (2019)
TriviaQA Exact Match 5 Joshi et al. (2017)
Code HumanEval Pass@1 0 Chen et al. (2021)
MBPP Pass@1 1 Austin et al. (2021)
Math GSM8K Exact Match 8 Cobbe et al. (2021)
General Ability MMLU Accuracy 5 Hendrycks et al. (2021)
BBH Accuracy 3 Suzgun et al. (2022)
@@ B.6 Model Customization
Code CodeFeedback 156k  Zheng et al. (2024a)
CodeAlpaca 20k Chaudhary (2023)
Math MathInstruction 262K Yue et al. (2023)
Medical MedMCQA 183k Zhang et al. (2023)
Law Legalbench 90k Guha et al. (2023)
Finance FinGPT 204k Wang et al. (2023a)
Alignment Ultrafeedback 62k Cui et al. (2024)
Customization Training Hyperparameters. In model customization, we use different hyperparameters depending on the model size. For LLaMA2-70B, we apply QLoRA with the settings outlined in Table 13, while for 7B models, we use LoRA. For smaller models like Phi2 and Phi-1.5, we fine-tune all model parameters. For LLaMA2-70B, we fine-tune it as a quantized 4-bit model over 1 epoch, starting with a learning rate of $1.5\times 10^{-6}$. For the 7B models, we train for 3 epochs, with a seed value of 42. The training setup includes a weight decay of 0.1, a batch size of 128, a warmup ratio of 0.03, and input sequences of 512 tokens, following standard experimental practices (Hu et al., 2021). For Phi2 and Phi-1.5, we use the training hyperparameters from the LLaMA2 report. We employ the AdamW optimizer with a cosine learning rate scheduler, starting with a learning rate of $2\times 10^{-5}$, a weight decay of 0.1, a batch size of 128, and use bfloat16 precision for 512-token input sequences. Specifically, for alignment, we follow SimPO Meng et al. (2024) and set the preference parameters $\beta=2$ and $\gamma=1$. The learning rate is $1\times 10^{-6}$ for LLaMA2-70B and $5\times 10^{-7}$ for the 7B and smaller models. All experiments are conducted using the LLaMA-Factory on Nvidia 4090 24G, 6000 Ada 48G, and A100 80G GPUs, with PyTorch 2.2.0 and CUDA 11.8 on Ubuntu 20.04.6 LTS.
@@
B.7 Amplification of small recovery error
To investigate the amplification of minor recovery errors in the pre-transition layers, we conducted experiments on the LLaMA2-7B model. Specifically, we added Gaussian noise, $x\sim N(0,0.01^{2})$, to the parameters of the first decoder layer of the model. Then, we compared the embedding outputs generated by the noisy model with those from the original model. The difference between the two sets of outputs was measured using the Frobenius norm. We collected embedding outputs from layers 0, 3, 6, 9, 12, 15, 18, 21, 24, 27, and 30, totaling 45,000 output samples. The results showed a significant amplification of these small errors.
B.8 Resilience and Customization Transitions
For the LLaMA2-7B model, the smallest closed-source layer set identified by SCARA consists of a single decoder layer, whereas for Phi-2, it includes two decoder layers. Consequently, for LLaMA2-7B, we opted to closed-source each even-indexed layer, while for Phi-2, we chose to closed-source non-overlapping pairs of layers (e.g., layers 0-1, 2-3). For each selected layer set, we first closed-source them, then subjected the semi-open model to FT-all attacks, and subsequently calculated the $\Delta\text{ARR}$ of the layer set to assess its resilience.
When verifying the customization transition, due to computational constraints, we validated only every other layer set for both models (e.g., closed-source layers 0, 0-4, 0-8 …). Specifically, we applied LoRA-based customization on LLaMA2-7B in the math domain, while for Phi-2, we utilized the full finetuning approach. The experimental hyperparameters remain consistent with those outlined in the Appendix B.6.
We further computed the $\Delta\text{ARR}$ for each closed-source set within Mistral-7B-v0.1 and Phi-1.5. In these models, the smallest closed-source set identified by SCARA consists of one decoder layer and two decoder layers, respectively. Following the same experimental configuration as LLaMA2-7B and Phi-2, we closed-sourced each even-indexed layer for Mistral-7B, and non-overlapping pairs of layers for Phi-1.5. The complete results demonstrating the transition layers within the Mistral-7B and Phi-1.5 model that closed two non-overlapping consecutive layers are depicted in Figure 10. Once again, we observed a distinct presence of transition layers. Specifically, in Mistral-7B, the transition layer appears at the 24th layer, while in Phi-1.5, it is located within the first layer set. Further results for can be found in Appendix C.7.
Figure 10: Resilience changes in Miatral-7B and Phi-1.5.
B.9 Resilience Across Closed Sizes
To examine the influence of Closed layer size on model resilience, we conduct experiments on Closed-sourcing different amounts and proportions of parameters in the model’s decoder layer. We give instructions on the detailed setting of closed-sourced models in Table 14. The module names are all derived from the overall implementation functions of each model in the Transformers open-source repositories in Table 6. We utilize abbreviated module names to denote specific settings.
Table 14: Closed-sourced Sizes Setting. “*” indicates an entire decoder layer.
@@ B.9 Resilience Across Closed Sizes
Proportion 100% Fully-closed Fully-closed Fully-closed Fully-closed
@@
We further computed $\Delta\text{ARR}$ by close-sourcing varying quantities and proportions of parameters under FT-all attacks on three additional models. As shown in Figure 11 and Figure 6(b), we observed the same pattern as with Llama2-7B, where resilience emerges once a sufficient number of parameters are closed-sourced. For example, on Mistral-7B, resilience occurs after closed-sourcing 100 million parameters, which is less than a single decoder layer. Closed-sourcing fewer parameters leads to a notable drop in resilience, with $\Delta\text{ARR}$ rising to around 40%. Beyond this threshold, resilience stabilizes near 0% $\Delta\text{ARR}$. This pattern holds across all models, highlighting a critical threshold for effective closed-source. Furthermore, different architectures require varying closed-sourcing quantities to achieve resilience, even with similar model sizes. For instance, Mistral-7B reaches resilience by closed-sourcing 100 million parameters, Llama2-7B requires 200 million, and Phi-1.5 needs a higher rate of 7%, compared to 3% for Llama2-7B.
Figure 11: $\Delta$ ARR for different closed parameter quantities and proportions.
B.10 Effectiveness of recovery difficulty
@@
The complete Pearson and Spearman results are presented in Table 15, revealing a negative correlation between RS and the average recovery ratio. For example, in Llama2-7B, both Pearson and Spearman coefficients fall below -0.77. Similar trends are seen in models with varying architectures and sizes, confirming that RD is a reliable predictor of recovered model performance and demonstrating the effectiveness of SCARA. Additionally, Figure 12 shows scatter plots depicting the relationship between $\Delta$ARR and Recovery Difficulty($\uparrow$)s across four models, along with the corresponding Pearson and Spearman correlation coefficients. The Recovery Difficulty($\uparrow$)s were obtained from Section 5.3. As illustrated in Figure 12, we observe a clear trend: an increase in $\Delta$ARR corresponds to a decrease in model scores across all models analyzed. This inverse relationship is consistently supported by strong negative values for both Pearson and Spearman correlation coefficients, with the most significant negative correlation seen in Phi2-2.7B, indicating a substantial drop in model scores as $\Delta$ARR increases.
@@
Table 15: Correlation coefficients (Spearman | Pearson) between recovery ratio and recovery difficult.
@@
Figure 12: Correlation Analysis of $\Delta$ARR and Recovery Difficulty Across Different Models.
B.11 Adversarial Attack
In this section, we provide a detailed comparison of SCARA and SAP-DP in their effectiveness against three types of black-box adversarial attacks on the Llama2-7B model. The attacks considered include Membership Inference Attacks (MIA), Attribute Inference Attacks (AIA), and Prompt Injection Attacks (PIA).
Membership Inference Attack (MIA): This attack aims to determine whether a specific data point was included in the training dataset of the model. Attackers utilize model outputs to infer membership status, potentially exposing sensitive information about the training data (Fu et al., 2023; Chen & Pattabiraman, 2024). We conducted our experiment following SPV_MIA  [^10] https://github.com/wjfu99/MIA-LLMs, which provides a robust framework for assessing model vulnerabilities. We focus on the AUC scores for SPV-MIA against semi-open models across Ag News datasets (Zhang et al., 2016).
Attribute Inference Attack (AIA): In this scenario, the adversary attempts to infer specific attributes of training data based on the model’s outputs. This can lead to privacy breaches, particularly when sensitive attributes are involved (Staab et al., 2023; Li et al., 2024). We conducted our experiments following the methodology outlined in Staab et al. (2023)  [^11] https://github.com/eth-sri/llmprivacy and evaluated the top-3 accuracy on the PersonalReddit (PR) Dataset.
Prompt Injection Attack (PIA): This attack manipulates input prompts to coerce the model into producing desired outputs that may compromise the integrity or security of the system (Zhao et al., 2024; Xu et al., 2024). In our experiment, we follow AutoDAN  [^12] https://github.com/SheltonLiu-N/AutoDAN, which can automatically generate stealthy jailbreak prompts by the carefully designed hierarchical genetic algorithm.We evaluate the effectiveness of these prompts using the keyword-based attack success rate (ASR), which measures the presence of predefined keywords in responses generated LLMs. For gold standard, LED  [^13] https://github.com/ledllm/ledllm, significantly enhances the resilience of LLMs against prompt injection attacks (PIA), reducing the ASR to 0.
@@ C.1 Comparison in two semi-open Llama2-70B
In this experiment, we examine two semi-open Llama2-70B models, where either the first two decoder layers are closed-source (referred to as Semi-Open-1) or the last two decoder layers are closed-source (referred to as Semi-Open-2). The objective is to compare their performance in terms of customization and their resilience under the recovery attack. The results are summarized in Table 16 and Table 17.
Table 16: Customization Performance of Llama2-70B under Different Closed-Sourced Layers
@@
Fully Closed-sourced 53.15 24.90 53.68 79.63 37.54 7.19
Semi-Open-1 62.40 43.99 62.73 93.85 87.51 7.46
Semi-Open-2 62.53 42.36 62.72 93.91 87.90 7.46
Table 17: Recovery Performance of Llama2-70B under Different Closed-Sourced Layers
Benchmarks Fully Closed-sourced Semi-Open-1 Semi-Open-2
@@
Average Recovery Ratio($\downarrow$) 22.55 21.73 74.94
@@ C.2 Evaluation Results under FT-all attack
In this section, we provide a comprehensive analysis of the evaluation results, comparing SCARA with two baseline methods: SAP-DP and a fully-closed approach. This comparison is conducted across 16 benchmarks under the FT-all attack scenario. The detailed results for Llama2-70B are presented in Table 18, while the results for Llama2-7B and Mistral-7B are shown in Table 19. Additionally, the outcomes for Phi-2 and Phi-1.5 are provided in Tables 20.
@@
Pre-train SCARA SAP-DP Fully-closed
@@
Average Recovery Ratio($\downarrow$) - 21.73 21.64 22.55
@@
SCARA SAP-DP Fully-closed SCARA SAP-DP Fully-closed
@@
Average Recovery Ratio($\downarrow$) 25.03 24.16 25.62 22.41 22.28 22.68
@@
SCARA SAP-DP Fully-closed SCARA SAP-DP Fully-closed

### v2

@@ B.5 Evaluation Benchmarks
Code HumanEval Pass@1 0 (Chen et al., 2021)
MBPP Pass@1 1 (Austin et al., 2021)
Math GSM8K Exact Match 8 (Cobbe et al., 2021)
Medical MMLU_Medical Accuracy 5 (Hendrycks et al., 2021)
Finance FPB F1 0 (Wang et al., 2023a)
Law LegalBench Accuracy 0 (Guha et al., 2023)
Alignment MT-Bench Score (GPT-4) (Zheng et al., 2023)
Evaluation on Security. We follow the Llama-2 report (Touvron et al., 2023) to evaluate the distilled model, including 16 benchmarks, which are categorized into 6 groups. Table 7 summarizes the functionality benchmarks used in our experiments, along with their test methods and performance metrics. Our model ranks choices in multiple-choice tasks and generates answers for open-ended generation tasks.
@@
Commonsense Reasoning PIQA Accuracy 0 (Bisk et al., 2020)
Hellaswag Accuracy 0 (Zellers et al., 2019)
Winogrande Accuracy 0 (Sakaguchi et al., 2019)
ARC_easy Accuracy 0 (Clark et al., 2018)
ARC_challenge Accuracy 0 (Clark et al., 2018)
Reading Comprehension OpenBookQ Accuracy 0 (Mihaylov et al., 2018)
LAMBADA Accuracy 0 (Paperno et al., 2016)
BoolQ Accuracy 0 (Clark et al., 2019)
SQuADv2 HasAns_EM 2 (Rajpurkar et al., 2018)
SQuADv2 HasAns_F1 2 (Rajpurkar et al., 2018)
World Knowledge NaturalQuestions Exact Match 5 (Kwiatkowski et al., 2019)
TriviaQA Exact Match 5 (Joshi et al., 2017)
Code HumanEval Pass@1 0 (Chen et al., 2021)
MBPP Pass@1 1 (Austin et al., 2021)
Math GSM8K Exact Match 8 (Cobbe et al., 2021)
General Ability MMLU Accuracy 5 (Hendrycks et al., 2021)
BBH Accuracy 3 (Suzgun et al., 2022)
@@ B.6 Model Customization
Code CodeFeedback 156k  (Zheng et al., 2024a)
CodeAlpaca 20k (Chaudhary, 2023)
Math MathInstruction 262K (Yue et al., 2023)
Medical MedMCQA 183k (Zhang et al., 2023)
Law Legalbench 90k (Guha et al., 2023)
Finance FinGPT 204k (Wang et al., 2023a)
Alignment Ultrafeedback 62k (Cui et al., 2024)
Customization Training Hyperparameters. In model customization, we use different hyperparameters depending on the model size. For LLaMA2-70B, we apply QLoRA with the settings outlined in Table 9, while for 7B models, we use LoRA. For smaller models like Phi2 and Phi-1.5, we fine-tune all model parameters. For LLaMA2-70B, we fine-tune it as a quantized 4-bit model over 1 epoch, starting with a learning rate of $1.5\times 10^{-6}$. For the 7B models, we train for 3 epochs, with a seed value of 42. The training setup includes a weight decay of 0.1, a batch size of 128, a warmup ratio of 0.03, and input sequences of 512 tokens, following standard experimental practices (Hu et al., 2021). For Phi2 and Phi-1.5, we use the training hyperparameters from the LLaMA2 report. We employ the AdamW optimizer with a cosine learning rate scheduler, starting with a learning rate of $2\times 10^{-5}$, a weight decay of 0.1, a batch size of 128, and use bfloat16 precision for 512-token input sequences. Specifically, for alignment, we follow SimPO (Meng et al., 2024) and set the preference parameters $\beta=2$ and $\gamma=1$. The learning rate is $1\times 10^{-6}$ for LLaMA2-70B and $5\times 10^{-7}$ for the 7B and smaller models. All experiments are conducted using the LLaMA-Factory on Nvidia 4090 24G, 6000 Ada 48G, and A100 80G GPUs, with PyTorch 2.2.0 and CUDA 11.8 on Ubuntu 20.04.6 LTS.
@@
B.7 Security and Customization Transitions
For the LLaMA2-7B model, the smallest secure-source layer set identified by SOLID consists of a single decoder layer, whereas for Phi-2, it includes two decoder layers. Consequently, for LLaMA2-7B, we opted to secure-source each even-indexed layer, while for Phi-2, we chose to secure-source non-overlapping pairs of layers (e.g., layers 0-1, 2-3). For each selected layer set, we first secure-source them, then subjected the semi-open model to FT-all attacks, and subsequently calculated the $\Delta\text{ADR}$ of the layer set to assess its security.
When verifying the customization transition, due to computational constraints, we validated only every other layer set for both models (e.g., secure-source layers 0, 0-4, 0-8 …). Specifically, we applied LoRA-based customization on LLaMA2-7B in the math domain, while for Phi-2, we utilized the full finetuning approach. The experimental hyperparameters remain consistent with those outlined in the Appendix B.6.
We further computed the $\Delta\text{ADR}$ for each secure-source set within Mistral-7B-v0.1 and Phi-1.5. In these models, the smallest secure-source set identified by SOLID consists of one decoder layer and two decoder layers, respectively. Following the same experimental configuration as LLaMA2-7B and Phi-2, we secured each even-indexed layer for Mistral-7B, and non-overlapping pairs of layers for Phi-1.5. The complete results demonstrating the transition layers within the Mistral-7B and Phi-1.5 model that secure two non-overlapping consecutive layers are depicted in Figure 7. Once again, we observed a distinct presence of transition layers. Specifically, in Mistral-7B, the transition layer appears at the 24th layer, while in Phi-1.5, it is located within the first layer set. Further results for can be found in Appendix C.7.
Figure 7: Security changes in Miatral-7B and Phi-1.5.
B.8 Security Across Secure Sizes
To examine the influence of Secure layer size on model security, we conduct experiments on Secure-sourcing different amounts and proportions of parameters in the model’s decoder layer. We give instructions on the detailed setting of secured models in Table 10. The module names are all derived from the overall implementation functions of each model in the Transformers open-source repositories in Table 3. We utilize abbreviated module names to denote specific settings.
Table 10: Secured Sizes Setting. “*” indicates an entire decoder layer.
@@ B.8 Security Across Secure Sizes
Proportion 100% Fully-secured Fully-secured Fully-secured Fully-secured
@@
We further computed $\Delta\text{ADR}$ by close-sourcing varying quantities and proportions of parameters under FT-all attacks on three additional models. As shown in Figure 8 and Figure LABEL:fig:appendix:appendix-specific-domain(a), we observed the same pattern as with Llama2-7B, where security emerges once a sufficient number of parameters are secured. For example, on Mistral-7B, security occurs after secure-sourcing 100 million parameters, which is less than a single decoder layer. Secure-sourcing fewer parameters leads to a notable drop in security, with $\Delta\text{ADR}$ rising to around 40%. Beyond this threshold, security stabilizes near 0% $\Delta\text{ADR}$. This pattern holds across all models, highlighting a critical threshold for effective secure-source. Furthermore, different architectures require varying secure-sourcing quantities to achieve security, even with similar model sizes. For instance, Mistral-7B reaches security by secure-sourcing 100 million parameters, Llama2-7B requires 200 million, and Phi-1.5 needs a higher rate of 7%, compared to 3% for Llama2-7B.
Figure 8: $\Delta$ ADR for different secure parameter quantities and proportions.
We explore how secured parameter ratio impacts the model security in Llama2-7B, as shown in Figure LABEL:fig:appendix:appendix-specific-domain(b). For instance, technical skills such as Math show earlier transitions, with security emerging at 1% parameters secured, whereas domains such as Commonsense Reasoning require hiding 3%. In summary, secure-sourcing a small portion of parameters can provide sufficient security against model distillation, meanwhile, technical capabilities tend to be more challenging to distill than other domains.
B.9 Effectiveness of distillation difficulty
@@
DD is effective in larger models, with a clear negative correlation between DD and average distillation ratios. As shown in Figure 9 (a), the Pearson coefficient for Llama2-7B consistently remains below -0.80, reaching as low as -0.98. Similar trends are observed in other models with varying architectures and sizes, confirming DD as a reliable predictor of distilled model performance and demonstrating the effectiveness of SOLID. Furthermore, the complete Pearson and Spearman results, presented in Table 11, reveal a negative correlation between RS and the average distillation ratio. For instance, in Llama2-7B, both the Pearson and Spearman coefficients fall below -0.77. These trends hold across models of different architectures and sizes. Additionally, Figure 10 displays scatter plots illustrating the relationship between $\Delta$ADR and Distillation Difficulty($\uparrow$)s for four models, with corresponding Pearson and Spearman correlation coefficients. The Distillation Difficulty($\uparrow$)s, obtained from Section 5.3, show a clear inverse relationship: an increase in $\Delta$ADR leads to a decrease in model scores across all models analyzed. This trend is consistently supported by strong negative correlations, with the most significant drop in model scores seen in Phi2-2.7B, where $\Delta$ADR shows the most substantial negative correlation.
DD is ineffective in smaller OPT model, with notable inconsistencies with ADRs. As shown in Figure 9 (b), the negative correlation between DD and ADR in OPT-350M is weak (coefficients above -0.33), showing its unsuitability for predicting distillation performance. Additionally, optimal security is achieved by protecting the middle layers rather than the initial or output layers, making SOLID unable to identify the smallest secured set.
Figure 9: (a) presents the Pearson coefficient between distillation ratio (DR) and distillation difficulty (DD) across four models and six domains. (b) depicts the link between ADR and DD for Llama2-7B and OPT-350M.
@@
Table 11: Correlation coefficients (Spearman | Pearson) between distillation ratio and distillation difficult.
@@
Figure 10: Correlation Analysis of $\Delta$ADR and Distillation Difficulty Across Different Models.
@@ C.1 Comparison in two semi-open Llama2-70B
In this experiment, we examine two semi-open Llama2-70B models, where either the first two decoder layers are secure-source (referred to as Bottom2-Secured) or the last two decoder layers are secure-source (referred to as Top2-Secured). The objective is to compare their performance in terms of customization and their security under the distillation attack. The results are summarized in Table 12 and Table 13.
Table 12: Customization Performance of Llama2-70B under Different Secured Layers
@@
Fully Secured 53.15 24.90 53.68 79.63 37.54 7.19
Bottom2-Secured 62.40 43.99 62.73 93.85 87.51 7.46
Top2-Secured 62.53 42.36 62.72 93.91 87.90 7.46
Table 13: Distillation Performance of Llama2-70B under Different Secured Layers
Benchmarks Fully Secured Semi-Open-1 Semi-Open-2
@@
Average Distillation Ratio($\downarrow$) 22.55 21.73 74.94
@@ C.2 Evaluation Results under FT-all attack
In this section, we provide a comprehensive analysis of the evaluation results, comparing SOLID with two baseline methods: SAP-DP and a fully-secured approach. This comparison is conducted across 16 benchmarks under the FT-all attack scenario. The detailed results for Llama2-70B are presented in Table 14, while the results for Llama2-7B and Mistral-7B are shown in Table 15. Additionally, the outcomes for Phi-2 and Phi-1.5 are provided in Tables 16.
@@
Pre-train SOLID SAP-DP Fully-secured
@@
Average Distillation Ratio($\downarrow$) - 21.73 21.64 22.55
@@
SOLID SAP-DP Fully-secured SOLID SAP-DP Fully-secured
@@
Average Distillation Ratio($\downarrow$) 25.03 24.16 25.62 22.41 22.28 22.68
@@
SOLID SAP-DP Fully-secured SOLID SAP-DP Fully-secured